### v2

@@ 1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2023b; 2024; Dubey et al., 2024; Deitke et al., 2024; OpenAI, 2024b; Team et al., 2023) have demonstrated impressive capabilities in tasks such as image captioning, visual question answering, and complex reasoning (Yue et al., 2024a; b). Despite this rapid progress in their reasoning capabilities, a critical flaw persists: the overwhelming focus on English- and western-centric training datasets and evaluation benchmarks (Liu et al., 2021; Song et al., 2023).
@@
In this paper, we address how to train and evaluate multilingual MLLMs that are culturally inclusive, using limited open-source resources, tackling four major challenges (Yu et al., 2022): 1) Data scarcity: high-quality multilingual multimodal data is scarce, especially in low-resource languages, which makes it difficult to create large-scale training data; 2) Cultural nuances: visual interpretations are context-dependent and vary across cultures (Ramaswamy et al., 2023; Khanuja et al., 2024); 3) Catastrophic forgetting: training on many languages or modalities often results in suboptimal performance on some subsets and require careful balancing; 4) Evaluation complexity: developing an evaluation suite that accurately measures performance across languages and cultures requires substantial resources and expertise.
To tackle these challenges, we introduce Pangea, an open-sourced multilingual MLLM designed to bridge linguistic and cultural gaps in visual understanding tasks. Pangea is trained on PangeaIns (Figure 2), a high-quality multilingual multimodal instruction tuning dataset comprising 6 million samples in 39 typologically diverse languages. PangeaIns combines existing open-source resources with newly created instructions focused on multicultural understanding. We curate high-quality English instructions, carefully translate them, and adapt them for multilingual contexts. To address Western-centric biases in visual representations, we source images from LAION-Multi (Schuhmann et al., 2022), which includes images from various countries and captions in multiple languages. However, LAION-Multi contains many images that are not culturally representative of the country’s speaking population. Additionally, the associated alt text is often short, noisy, and lacks sufficient detail. To combat these issues, we develop a multi-cultural multilingual multimodal instruction generation pipeline. This pipeline leverages an LLM (Dubey et al., 2024) to score and filter images based on cultural informativeness. We then enhance the remaining data by generating detailed descriptions and creating complex instructions that combine culturally relevant tasks with general multilingual scenarios. This approach improves the model’s cultural understanding while maintaining robust multilingual performance.
@@
Our results demonstrate Pangea’s abilities in both English and multilingual scenarios, significantly outperforming existing open-source MLLMs across PangeaBench datasets, surpassing the best open MLLMs by 0.4 points on English tasks and 10.9 points on multilingual tasks on average. Notably, Pangea excels in multilingual and multi-cultural understanding, as evidenced by its performance on xChat, CVQA, and MaRVL benchmarks. Pangea also matches or outperforms state-of-the art proprietary LLMs, namely Gemini-1.5-Pro and GPT4o, on several tasks such as XGQA. However, some performance gaps remain in multimodal chat and complex reasoning, shedding light on the need for further improvements in open multimodal LLMs. We discuss key insights from Pangea’s training, including the scaling effect of instructions, the role of English data, and the impact of language-specific training proportions on performance. Additionally, we explore preliminary methods to improve multilingual OCR. We fully open-source PangeaIns, PangeaBench, Pangea models, and code, to advance culturally inclusive MLLMs across diverse languages.
Figure 2: Statistics of our PangeaIns, which comprises 6M multimodal instructions in 39 languages. The distribution of multilingual training samples shows the percentage of instances for each language among the multilingual instructions. PangeaIns includes general instructions, document and chart question answering, captioning, domain-specific, culturally relevant, and text-only instructions. To address the data scarcity of multilingual multimodal instructions, we employ three strategies: 1) translating high-quality English instructions into multiple languages, filtering culturally relevant images; 2) generating culturally aware instructions, and 3) collecting and re-purposing existing open-source multimodal datasets.
@@ 2 PangeaIns: Multilingual Multimodal Instruction Tuning
Creating a truly multilingual, multicultural MLLM presents unique challenges. We developed PangeaIns, a diverse and high-quality dataset specifically designed for instruction tuning. Comprising 6 million samples in 39 languages, PangeaIns was curated with a focus on linguistic and cultural diversity. We implemented three key strategies to ensure comprehensive coverage, each addressing the specific hurdles encountered in multilingual multimodal learning. Figure 2 shows the details of our PangeaIns’s distribution.
@@ 2.1 Machine Translated Instructions
The scarcity of human-annotated multilingual multimodal training datasets presents a significant challenge. To address this, we primarily adopt machine translation as a practical and scalable solution to extend our dataset beyond English. While human annotation is ideal, it is resource-intensive and impractical for covering a wide range of languages.
@@ 2.2 Multicultural Understanding Instructions
Captioning of Multicultural Images with Different Languages. To provide context and enhance the models’ ability to interpret the images accurately, we regenerated a more detailed caption using Gemini 1.5 Pro based on high-quality original text. Each image was accompanied by a caption written in the language corresponding to its cultural origin. This step was crucial as writing captions in the native language of the image’s cultural context adds a layer of authenticity and helps the model learn language-specific nuances.
Generating Multilingual and Cross-Cultural Instructions. For each image, we generated captions in the corresponding native language using Gemini 1.5 Pro. However, our approach was not just about using a capable model. The filtered high-quality alt text played a critical role in enriching the data, as it often contained culturally specific and contextually important information that would otherwise be absent from the images alone. For example, in the Appendix Figure 8, with high-quality alt text, a model can incorporate details such as “President and CEO of The Walt Disney Company standing in front of a model of Shanghai Disneyland,” which adds significant context that may not be immediately evident from the image. This additional layer of information helps the model generate captions that better capture the cultural and contextual nuances.
@@ 2.3 Curating Existing Multilingual Instructions
To further enrich PangeaIns, we conducted an extensive survey of the available multilingual multimodal literature and datasets, including those hosted on HuggingFace. As a result, we incorporated several high-quality, open-source datasets into our PangeaIns mixture. These include Chinese ALLaVA-4V (Chen et al., 2024), Viet Document and OCR QA (Doan et al., 2024), Llava Chinese (LinkSoul-AI, 2023), Llava Medical Chinese Instruction (BUAA, 2023), LLaVA-Japanese-Instruct (Toshi456, 2023), MTVQA (Tang et al., 2024), Japanese STAIR Captions (Yoshikawa et al., 2017), Russian GQA (Belopolskih & Spirin, 2024), French Doc-VQA (Sonagu & Sola, 2024), and French Table-VQA (Agonnoude & Delestre, 2024). Each of these datasets brings unique linguistic and cultural perspectives to the mix, covering a wide range of languages and task types.
@@ 2.4 Dataset Statistics
By combining these three methods, we created PangeaIns as a comprehensive dataset that addresses the major challenges in building multilingual multimodal models: data scarcity, linguistic diversity, and cultural nuance. The dataset’s balanced distribution across languages and tasks supports the development of more sophisticated LLMs that can handle complex visual and textual content in a multilingual, multicultural context.
Language and Task Distribution: PangeaIns features an extensive and balanced distribution of languages, tasks, and cultural contexts (as shown in Figure 2). We empirically keep the final language ratio of English to Multilingual as 40%:60% as we found a significant portion of English data plays an important role in cross-lingual transfer. See more discussions about the ratio in section 5 and Figure 6. The inclusion of diverse multimodal instructions ensures that the model develops a deeper understanding of varied linguistic and cultural environments. Examples of training samples from different languages and categories are provided in Appendix H.
The comprehensive nature of PangeaIns lays a solid foundation for training Pangea, enabling it to become a truly multilingual, multicultural multimodal LLM, capable of understanding and interacting effectively with users from diverse linguistic and cultural backgrounds.
@@ 3.2 Multimodal Tasks
The multimodal tasks in PangeaBench are categorized as follows: Multimodal Chat, Captioning, Cultural Understanding, Multilingual Visual Question Answering (VQA), and Multi-Subject Reasoning. We incorporate these tasks in PangeaBench to ensure comprehensive testing of Pangea’s multimodal capabilities. The overview and examples of PangeaBench from each task are shown in Figure 4.
Multimodal Chat. This task tests the model’s ability to engage in natural and dynamic real-world conversations involving both text and images. In this category, Multilingual LlavaBench (Rasheed et al., 2025) (M-LlavaBench for short) stands as the only benchmark for evaluating multilingual long-form generation capabilities from MLLMs. Moreover, following the evaluation pipeline from Zheng et al. (2023) and Liu et al. (2023b), M-LlavaBench uses a coarse-grained evaluation criteria (for example, “Please rate the helpfulness, relevance, accuracy, level of details of their responses.”). Previous works have shown that employing such coarse-grained evaluation criteria might lead to automatic evaluation results that diverge from how humans would evaluate them (Ye et al., 2023; Kim et al., 2023; Lee et al., 2024; Kim et al., 2024a; b). To assess our baselines based on a more accurate evaluation pipeline with fine-grained evaluation criteria on diverse scenarios, we additionally annotate a new multilingual multimodal generation benchmark called the xChatBench, included in the multimodal chat category of PangeaBench. A more detailed explanation of the annotation process of xChatBench is included in Appendix E.
Captioning. The XM3600 (Thapliyal et al., 2022) dataset was developed to evaluate models’ capability in multilingual image captioning. This dataset contains images paired with captions in 36 different languages. The original datasets often contain similar images and captions. To address this, we clustered the images based on their captions and manually selected 100 representative images from these clusters. This approach enhances the diversity of the dataset while also accelerating the evaluation process. We call the dataset XM100.
@@
Multi-Subject Reasoning. The xMMMU and M3Exam (Zhang et al., 2023) datasets are used to evaluate the model’s reasoning abilities across different academic subjects. xMMMU is a machine-translated version of MMMU validation questions, which focuses on multimodal reasoning in multiple subjects. We randomly sample 300 questions from MMMU (Yue et al., 2024a) validation set and employ GPT-4o for the six languages translation. M3Exam challenges the model with complex, real-world educational questions requiring both textual and visual comprehension. Further details on how we ensure the quality of translations for xMMMU, as well as a detailed description of other datasets, can be found in Appendix D.
@@ 3.3 Text-Only Multilingual Datasets
Specifically, we include TydiQA (Clark et al., 2020) to test the model’s ability to answer questions across 11 typologically diverse languages. We adopt the FLORES (NLLB Team, 2024) dataset to assess machine translation performance across multiple languages. We sample 11 languages (denoted as FLORES-Sub). We use MMMLU (OpenAI, 2024a), a human-translated version of MMLU to test the general language understanding of models. We use XStoryCloze (Lin et al., 2021) and MGSM (Shi et al., 2022) to test the model’s commonsense and mathematical reasoning ability in multilingual contexts respectively.
@@ 4.1 Experimental Setup
We train Pangea on PangeaIns, our multilingual multimodal dataset comprising 6 million samples across 39 languages. The model is based on LLaVA-Next architecture (Liu et al., 2024) with Qwen2-7B-Instruct (Yang et al., 2024) as the language model backbone and clip-vit-large-patch14-336 (Radford et al., 2021) as the vision encoder. The training consists of two stages. First, we pretrain the vision-language connector that aligns the outputs of clip-vit-large-patch14-336 to Qwen2-7B-Instruct, with the LLaVA LCS-558K [^1] https://huggingface.co/datasets/liuhaotian/LLaVA-Pretrain (Liu et al., 2023b; a). Then, we perform finetuning on PangeaIns, where we employ a learning rate of 2e-5, a batch size of 512, coupled with a cosine decay schedule with 0.03 warmup steps. We pretrain and finetune the model for 1 epoch, where pretraining took 4 hours with 8 H100 (32 GPU hours), and finetuning took 168 hours with 8 H100 (1344 GPU hours).
@@ 5 Discussion
Finally, we explore the implications of our findings and their potential impact on future developments in the field. We examine the scaling effects of instruction quantity, the persistent role of English data, the relationship between training sample proportions and performance, qualitative examples of model behavior, and challenges in multilingual OCR. Through this discussion, we aim to provide a comprehensive understanding of our model and chart a course for future advancements.
@@
As shown in Figure 6 , English performance generally improves as the percentage of English data increases. More surprisingly, using no English data (full multilingual data) results in relatively lower multilingual performance. As we introduce more English data, multilingual performance improves, peaking at 38.7% with 40% English. However, performance drops sharply when English data reaches 100%. This suggests that English data aids cross-lingual transfer, however, over-reliance on it harms multilingual performance.
How does the proportion of training samples in a language affect downstream performance? An interesting question to ask is whether the downstream task performance is correlated with the number of training samples. Our analysis in Figure 7 revealed a nuanced relationship between training sample proportion and downstream performance. While there is a general positive correlation, the impact varies significantly across languages and tasks. For widely spoken languages with rich linguistic resources, we observed a near-linear relationship. However, for low-resource languages, even a small increase in proportion yielded disproportionately large performance gains. Interestingly, we also noted instances of positive transfer between typologically similar languages. These findings suggest that strategic allocation of training samples, considering both language prevalence and linguistic similarities, can optimize overall model performance.
@@
Multimodal Chat in the Wild. One important application of MLLMs is to answer users’ queries in the wild. Here, we show the outputs of Pangea for the multimodal chat queries from our xChatBench. As shown in Appendix Figure 9, 10, 11, 12, 13, 14, Pangea successfully interprets the figures in different tasks and generates fluent and readable in certain languages. These qualitative examples further demonstrate the remarkable visual understanding ability of Pangea in multilingual contexts. On the other hand, we also identified a few bad cases shown in Figure 15, 16. Despite generating relevant responses to the queries, the model does not capture the key details of the images due to the lack of knowledge, which points out potential improvement directions in the future.
Preliminary Explorations of Multilingual OCR. Multilingual OCR emerged as a particularly challenging aspect of Pangea’s functionality. We made efforts to enhance its multilingual OCR capabilities. Further details on exploring multilingual OCR are available in Appendix J.

### v3

@@ 1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2023b; 2024; Dubey et al., 2024; Deitke et al., 2024; OpenAI, 2024b; Team et al., 2023) have demonstrated impressive capabilities in tasks such as image captioning, visual question answering, and complex reasoning (Yue et al., 2024a; b). Despite this rapid progress in their reasoning ability, a critical flaw persists: the overwhelming focus on English- and western-centric training and evaluation datasets (Liu et al., 2021; Song et al., 2023).
@@
In this paper, we address how to train and evaluate culturally inclusive multilingual MLLMs, using limited open-source resources, tackling four major challenges (Yu et al., 2022): 1) Data scarcity: high-quality multilingual multimodal data is scarce, especially in low-resource languages, making it difficult to create large-scale training data; 2) Cultural nuances: visual interpretations are context-dependent and vary across cultures (Ramaswamy et al., 2023; Khanuja et al., 2024); 3) Catastrophic forgetting: training on many languages or modalities often results in suboptimal performance on some subsets and require careful balancing; 4) Evaluation complexity: substantial resources and expertise are required to accurately measure performance across languages and cultures.
To tackle these challenges, we introduce Pangea, an open-source multilingual MLLM designed to bridge linguistic and cultural gaps in visual understanding tasks. Pangea is trained on PangeaIns (Figure 2), a high-quality multilingual multimodal instruction tuning dataset comprising 6 million samples in 39 typologically diverse languages. PangeaIns combines existing open-source resources with newly created instructions focused on multicultural understanding. We curate high-quality English instructions, carefully translate and adapt them for multilingual contexts. To address Western-centric biases in visual representations, we source images from LAION-Multi (Schuhmann et al., 2022), which includes images from various countries and captions in multiple languages. However, LAION-Multi contains images that are not culturally representative of the country’s speaking population, and the associated alt text is often short, noisy, and lacks sufficient detail. To combat these issues, we develop a multicultural multilingual multimodal instruction generation pipeline, leveraging an LLM (Dubey et al., 2024) to score and filter images based on cultural informativeness. We then enhance the remaining data by generating detailed descriptions and creating complex instructions that combine culturally relevant tasks with general multilingual scenarios. This approach improves the model’s cultural understanding while maintaining robust multilingual performance.
@@
Our results demonstrate Pangea’s abilities in both English and multilingual scenarios, significantly outperforming existing open-source MLLMs on PangeaBench, surpassing the best open MLLMs by 0.4% on English tasks and 10.9% on multilingual tasks on average. Notably, Pangea excels in multilingual and multicultural understanding, evidenced by its performance on xChat, CVQA, and MaRVL benchmarks. Pangea also matches or outperforms state-of-the-art proprietary LLMs, namely Gemini-1.5-Pro and GPT4o, on several tasks such as XGQA. However, some performance gaps remain in multimodal chat and complex reasoning, shedding light on the need for further improvements in open MLLMs. We discuss key insights from trainingPangea, including the scaling effect of instructions, the role of English data, the impact of language-specific training proportions, and preliminary methods to improve multilingual OCR. We fully open-source PangeaIns, PangeaBench, Pangea-7B, and code, to advance culturally inclusive MLLMs across diverse languages.
Figure 2: Statistics of PangeaIns, comprising 6M multimodal instructions in 39 languages. The distribution of multilingual training data shows the percent of instances for each language among the multilingual instructions. PangeaIns includes general instructions, document and chart question answering, captioning, domain-specific, culturally relevant, and text-only instructions.
@@ 2 PangeaIns: Multilingual Multimodal Instruction Tuning
Creating a truly multilingual, multicultural MLLM presents unique challenges. We developed PangeaIns, a diverse and high-quality instruction tuning dataset. Comprising 6 million samples in 39 languages, PangeaIns was curated with a focus on linguistic and cultural diversity. We implemented three key strategies to ensure comprehensive coverage, each addressing the specific hurdles encountered in multilingual multimodal learning. Figure 2 shows the distribution of PangeaIns.
@@ 2.1 Machine Translated Instructions
To address the scarcity of human-annotated multilingual multimodal data, we primarily adopt machine translation as a practical and scalable solution to extend data beyond English. While human annotation is ideal, it is resource-intensive and impractical to cover a wide range of languages.
@@ 2.2 Multicultural Understanding Instructions
Captioning Images with Different Languages. To provide context and enhance the model’s ability to interpret images, we regenerated more detailed captions using Gemini 1.5 Pro based on high-quality alt texts. In this step, each image was accompanied by a caption written in the language corresponding to its cultural origin. However, our approach was not just about using a capable model. The alt text played a critical role in enriching the data, as it often contained culturally specific and contextually important information that would otherwise be absent from the images alone. For example, in Figure 8, with high-quality alt text, models can incorporate details such as “President and CEO of The Walt Disney Company” and “a model of Shanghai Disneyland,” adding significant context that may not be immediately evident from the image. This additional layer of information helps the model generate captions that better capture the cultural and contextual nuances.
Generating Multilingual and Cross-Cultural Instructions.
@@ 2.3 Curating Existing Multilingual Instructions
To further enrich PangeaIns, we conducted an extensive survey of available multilingual multimodal literature and datasets, including those hosted on HuggingFace. As a result, we incorporated several high-quality, open-source datasets into PangeaIns. These include Chinese ALLaVA-4V (Chen et al., 2024), Viet Document and OCR QA (Doan et al., 2024), Llava Chinese (LinkSoul-AI, 2023), Llava Medical Chinese Instruction (BUAA, 2023), LLaVA-Japanese-Instruct (Toshi456, 2023), MTVQA (Tang et al., 2024), Japanese STAIR Captions (Yoshikawa et al., 2017), Russian GQA (Belopolskih & Spirin, 2024), French Doc-VQA (Sonagu & Sola, 2024), and French Table-VQA (Agonnoude & Delestre, 2024). Each of these datasets brings unique linguistic and cultural perspectives to the mix, covering a wide range of languages and task types.
@@ 2.4 Dataset Statistics
By combining these three methods, we created PangeaIns, a comprehensive dataset addressing major challenges in building multilingual MLLMs: data scarcity, linguistic diversity, and cultural nuance. Its balanced language and task distribution supports the development of more sophisticated LLMs that can handle complex visual and textual content in a multilingual, multicultural context.
Language and Task Distribution: PangeaIns features an extensive and balanced distribution of languages, tasks, and cultural contexts (as shown in Figure 2). We empirically keep the final language ratio of English to Multilingual as 40%:60% as we found a significant portion of English data plays an important role in cross-lingual transfer. See more discussions about the ratio in section 5 and Figure 6. The inclusion of diverse multimodal instructions ensures that the model develops a deeper understanding of varied linguistic and cultural environments. Examples of training samples from different languages and categories are provided in Appendix H. The comprehensive nature of PangeaIns lays a solid foundation for training Pangea, enabling it to become a truly multilingual, multicultural multimodal LLM, capable of understanding and interacting effectively with users from diverse linguistic and cultural backgrounds.
@@ 3.2 Multimodal Tasks
The multimodal tasks in PangeaBench are categorized as follows: Multimodal Chat, Captioning, Cultural Understanding, Multilingual Visual Question Answering (VQA), and Multi-Subject Reasoning. We incorporate these in PangeaBench to ensure comprehensive testing of Pangea’s multimodal capabilities. The overview and examples of PangeaBench are shown in Figure 4.
Multimodal Chat. This task tests the model’s ability to engage in natural and dynamic real-world conversations involving both text and images. Multilingual LlavaBench (Rasheed et al., 2025) (M-LlavaBench for short) stands as the only benchmark for evaluating multilingual long-form generation capabilities from MLLMs. Following the evaluation pipeline from Zheng et al. (2023) and Liu et al. (2023b), M-LlavaBench uses a coarse-grained evaluation criteria (e.g., “Please rate the helpfulness, relevance, accuracy, level of details of their responses.”). Previous works suggest that employing such coarse-grained evaluation criteria may lead to automatic evaluation results that diverge from how humans would evaluate them (Ye et al., 2023; Kim et al., 2023; Lee et al., 2024; Kim et al., 2024a; b). To assess baselines with a more accurate evaluation pipeline with fine-grained evaluation criteria on diverse scenarios, we additionally annotate a new multilingual multimodal generation benchmark called the xChatBench, included in the multimodal chat category of PangeaBench. A more detailed explanation of the annotation process of xChatBench is included in Appendix E.
Captioning. The XM3600 (Thapliyal et al., 2022) dataset was developed to evaluate models’ capability in multilingual image captioning. It contains images paired with captions in 36 different languages. However, it includes many similar images and captions. To address this, we clustered the images based on captions and manually selected 100 representative images (denoted as XM100). This approach enhances the diversity of the samples and accelerates the evaluation process.
@@
Multi-Subject Reasoning. The xMMMU and M3Exam (Zhang et al., 2023) datasets are used to evaluate the model’s reasoning abilities across different academic subjects. xMMMU is a machine-translated version of MMMU validation questions, which focuses on multimodal reasoning in multiple subjects. We randomly sample 300 questions from MMMU (Yue et al., 2024a) validation set and employ GPT-4o for the six languages translation. M3Exam challenges the model with real-world educational questions requiring both textual and visual comprehension. Details on how we ensure the translation quality, as well as detailed descriptions of other datasets, can be found in Appendix D.
@@ 3.3 Text-Only Multilingual Datasets
Specifically, we include TydiQA (Clark et al., 2020) to test the model’s ability to answer questions across 11 typologically diverse languages. We adopt the FLORES (NLLB Team, 2024) dataset to assess machine translation performance. We sample 11 languages (denoted as FLORES-Sub). We use MMMLU (OpenAI, 2024a), a human-translated version of MMLU to test the general language understanding. We use XStoryCloze (Lin et al., 2021) and MGSM (Shi et al., 2022) to test the model’s commonsense and mathematical reasoning ability in multilingual contexts respectively.
@@ 4.1 Experimental Setup
We train Pangea on PangeaIns, our multilingual multimodal dataset comprising 6 million samples across 39 languages. The model uses LLaVA-Next as architecture (Liu et al., 2024), Qwen2-7B-Instruct (Yang et al., 2024) as the language model backbone and clip-vit-large-patch14-336 (Radford et al., 2021) as the vision encoder. The training consists of two stages. First, we pretrain the vision-language connector that aligns the outputs of vision encoder to backbone, with the LLaVA LCS-558K [^1] https://huggingface.co/datasets/liuhaotian/LLaVA-Pretrain (Liu et al., 2023b; a). Then, we perform finetuning on PangeaIns, where we employ a learning rate of 2e-5, a batch size of 512, coupled with a cosine decay schedule with 0.03 warmup steps. We pretrain and finetune the model for 1 epoch, where pretraining took 4 hours with 8 H100 (32 GPU hours), and finetuning took 168 hours with 8 H100 (1344 GPU hours).
@@ 5 Discussion
Finally, we explore implications of our findings and their potential impact on future developments in the field. We examine the scaling effects of instruction quantity, the persistent role of English data, the relationship between training sample proportions and performance. Through this discussion, we aim to provide a comprehensive understanding of our model and chart a course for future advancements. More discussion on qualitative examples of model behavior on multilingual multimodal chat and challenges in multilingual OCR can be found in Appendix E and Appendix J.
@@
As shown in Figure 6 , English performance generally improves as the percent of English data increases. Surprisingly, using only multilingual data results in relatively lower multilingual performance. As we introduce more English data, multilingual performance improves, peaking at 38.7% with 40% English. However, performance drops sharply when English data reaches 100%. This suggests that English data aids cross-lingual transfer, however, over-reliance on it harms multilingual performance.
How does the proportion of training samples in a language affect downstream performance? Is downstream task performance correlated with the number of training samples? Our analysis in Figure 7 revealed the relationship between training sample proportion and downstream performance. While there is a general positive correlation, the impact varies significantly across languages and tasks. For widely spoken languages with rich resources, we observed a near-linear relationship. However, for low-resource languages, even a small increase in proportion yielded disproportionately large performance gains. Interestingly, we also noted instances of positive transfer between typologically similar languages. These findings suggest that strategic allocation of training samples, considering both language prevalence and linguistic similarities, can optimize model performance.